# Dynamic Subgraph Matching via Cost-Model-based Vertex Dominance Embeddings (Technical Report)

### v2

@@ 7. Cost-Model-Guided Embedding Optimization
In this section, we will design a novel cost model for evaluating the pruning power of our proposed vertex dominance embedding technique. Then, we will utilize this cost model to guide the embeddings and achieve better performance of dynamic subgraph matching.
7.1. Cost Model for Dynamic Subgraph Matching via Vertex Embeddings
In this subsection, we will propose a cost model to evaluate the performance of our dynamic subgraph matching (or the pruning power of using vertex dominance embeddings).
Specifically, we estimate the number, $Cost_{DSM}$, of candidate vertices $v_{i}$ (to retrieve and refine) whose embedding vectors $o^{\prime}(v_{i})$ are dominated by query embedding vector $o^{\prime}(q_{i})$ (given by Lemma 6.1). That is, we have:
(7) $Cost_{DSM}=\sum_{v_{i}\in V(G_{t})}\prod_{j=1}^{2d}Pr\{o^{\prime}(q_{i})[j]\leq o^{\prime}(v_{i})[j]\},$
@@ 7.1. Cost Model for Dynamic Subgraph Matching via Vertex Embeddings
(8) $Pr\{o^{\prime}(q_{i})[j]\leq o^{\prime}(v_{i})[j]\}$
$=$ $Pr\left\{\frac{(o^{\prime}(q_{i})[j]-o^{\prime}(v_{i})[j])-(o^{\prime}(q_{i})[j]-\mu_{o^{\prime}(v_{i})[j]})}{\sigma_{o^{\prime}(v_{i})[j]}^{2}}\right.$
$\left.\leq\frac{-(o^{\prime}(q_{i})[j]-\mu_{o^{\prime}(v_{i})[j]})}{\sigma_{o^{\prime}(v_{i})[j]}^{2}}\right\}.$
@@
(10) $Cost_{DSM}=|V(G_{t})|\cdot\prod_{j=1}^{2d}\Phi\left(\frac{\mu_{o^{\prime}(v_{i})[j]}-o^{\prime}(q_{i})[j]}{\sigma_{o^{\prime}(v_{i})[j]}^{2}}\right).$
7.2. Design of Cost-Model-Based Vertex Dominance Embeddings
Eq. (10) estimates the (worst-case) query cost for one query vertex $q_{i}$ during the dynamic subgraph matching. Intuitively, our goal is to design vertex embeddings $o^{\prime}(\cdot)$ that minimizes the cost model $Cost_{DSM}$ (given in Eq. (10)).
Note that, $\Phi(\cdot)$ in Eq. (10) is a monotonically increasing function. Therefore, if we can minimize the term $\frac{\mu_{o^{\prime}(v_{i})[j]}-o^{\prime}(q_{i})[j]}{\sigma_{o^{\prime}(v_{i})[j]}^{2}}$ in Eq. (10), then low query cost $Cost_{DSM}$ can be achieved. In other words, guided by our proposed cost model (to minimize the query cost $Cost_{DSM}$), our target is to design/select a “good” distribution of vertex dominance embeddings $o^{\prime}(v_{i})$ with:
@@ 7.2. Design of Cost-Model-Based Vertex Dominance Embeddings
In this paper, unlike standard vertex dominance embeddings generated by Uniform random function $f_{1}(\cdot)$ (as discussed in Section 4.2), we choose to use a (seeded) Zipf random function to produce SPUR vectors $x_{i}$ in vertex embeddings $o^{\prime}(v_{i})$ (or in turn generate SPAN vectors $y_{i}=\sum x_{j}$). The $Zipf$ distribution exactly follows our target of finding a random variable with low mean and high variance, which can achieve low query cost, as guided by our proposed cost model $Cost_{DSM}$ (given in Eq. (7)). We would like to leave the interesting topic of studying other low-mean/high-variance distributions as our future work.
@@
8. Experimental evaluation
8.1. Experimental Settings
To evaluate the performance of our DSM approach, we conduct experiments on a Ubuntu server equipped with an Intel Core i9-12900K CPU and 128GB memory. Our source code in C++ and real/synthetic graph data sets are available at URL: https://github.com/JamesWhiteSnow/DSM.
Real/Synthetic Graph Data Sets: We evaluate our DSM approach over both real and synthetic graphs.
Real-world graphs. We test five real-world graph data used by previous works (He and Singh, 2008; Shang et al., 2008; Zhao and Han, 2010; Lee et al., 2012; Sun et al., 2012; Han et al., 2013; Ren and Wang, 2015; Bi et al., 2016; Katsarou et al., 2017; Bhattarai et al., 2019; Han et al., 2019; Sun and Luo, 2020), which can be classified into three categories: i) biology networks (Yeast and HPRD); ii) bibliographical/social networks (DBLP and Youtube), and; iii) citation networks (US Patents). Statistics of these real graphs are summarized in Table 2.
@@ 8.1. Experimental Settings
Table 3. Parameter settings.
Parameters Values
the dimension, $d$, of the SPUR/SPAN vector 1 2, 3, 4
the ratio, $\beta/\alpha$ 10, 100, 1,000, 10,000, 100,000
the number, $m$, of degree groups 1, 2, 3, 4, 5
the number, $K$, of cell intervals in each dimension of $Syn_{j}$ 1, 2, 5, 8, 10
the number, $|\Sigma|$, of distinct labels 5, 10, 15, 20, 25
the average degree, $avg\_deg(q)$, of the query graph $q$ 2, 3, 4
the size, $|V(q)|$, of the query graph $q$ 5, 6, 8, 10, 12
the average degree, $avg\_deg(G_{D})$, of dynamic graph $G_{D}$ 3, 4, 5, 6, 7
the size, $|V(G_{D})|$, of the dynamic data graph $G_{D}$ 10K, 30K, 50K, 80K, 100K, 500K, 1M
Synthetic graphs. We generate synthetic graphs via NetworkX (Hagberg and Conway, 2020), and produce small-world graphs following the Newman-Watts-Strogatz model (Watts and Strogatz, 1998). Parameter settings of synthetic graphs are depicted in Table 3. For each vertex $v_{i}$, we generate its label $l(v_{i})$ by randomly picking up an integer within $[1,|\sum|]$, following the $Uniform$, $Gaussian$, or $Zipf$ distribution. Accordingly, we obtain 3 types of graphs, denoted as $Syn\text{-}Uni$, $Syn\text{-}Gau$, and $Syn\text{-}Zipf$, resp.
The insertion (or deletion) rate is defined as the ratio of the number of edge insertions (or deletions) to the total number of edges in the raw graph data. Following the literature (Choudhury et al., 2015; Kim et al., 2018; Min et al., 2021; Sun et al., 2022a), we set the insertion rate to 10% by default (i.e., the initial graph $G_{0}$ contains 90% edges, and the remaining 10% edges arrive as the insertion stream). Similarly, for the case of edge deletions, we set the default deletion rate to 10% (i.e., initial graph $G_{0}$ contains all edges, and deletion operations of 10% edges arrive in the stream).
Query Graphs: Similar to previous works (Sun et al., 2012; Han et al., 2013; Ren and Wang, 2015; Bi et al., 2016; Katsarou et al., 2017; Archibald et al., 2019; Bhattarai et al., 2019; Han et al., 2019; Sun et al., 2022a), for each graph $G_{D}$, we randomly extract/sample 100 connected subgraphs as query graphs, where parameters of query graphs $q$ (e.g., $|V(q)|$ and $avg\_deg(q)$) are depicted in Table 3.
Baseline Methods: We compare our DSM approaches (using the cost-model-based vertex dominance embeddings $o^{\prime}_{C}(v_{i})$, as discussed in Section 7) with five representative baseline methods of dynamic subgraph matching as follows:
@@
We used the code of baseline methods from (Sun et al., 2022a), which is implemented in C++ for a fair comparison.
(a) DSM vs. $d$
(b) DSM vs. $\beta/\alpha$
(c) DSM vs. $m$
(d) DSM vs. $K$
(e) DSM vs. embeddings
Figure 8. The DSM efficiency w.r.t parameters $\bm{d}$, $\bm{\beta/\alpha}$, $\bm{m}$, $\bm{K}$, and embedding design strategies.
Evaluation Metrics: In our experiments, we report the efficiency of our DSM approach and baseline methods, in terms of the wall clock time, which is the time cost of processing the DSM query over the entire graph update sequence (i.e., continuous subgraph matching answer set monitoring). In particular, this time cost includes the time of filtering, refinement, embedding updates, and synopsis updates for each DSM query. We also evaluate the pruning power of our embedding dominance pruning and MBR range pruning strategies (as mentioned in Section 6.1), which is defined as the percentage of vertices that can be ruled out by our pruning methods. For all the experiments, we take an average of each metric over 100 runs (w.r.t. 100 query graphs, resp.).
Table 3 depicts parameter settings in our experiments, where default parameter values are in bold. For each set of experiments, we vary the value of one parameter while setting other parameters to their default values.
(a) real-world graphs
(b) synthetic graphs
Figure 9. The DSM pruning power w.r.t different design strategies of vertex dominance embeddings.
(a) real-world graphs (10%)
(b) real-world graphs (20%)
(c) real-world graphs (30%)
(d) real-world graphs (40%)
(e) real-world graphs (50%)
(f) synthetic graphs (10%)
(g) synthetic graphs (20%)
(h) synthetic graphs (30%)
(i) synthetic graphs (40%)
(j) synthetic graphs (50%)
Figure 10. The DSM efficiency on edge-insertion-only real/synthetic graphs, compared with baseline methods.
(a) real-world graphs (10%)
(b) real-world graphs (20%)
(c) real-world graphs (30%)
(d) real-world graphs (40%)
(e) real-world graphs (50%)
(f) synthetic graphs (10%)
(g) synthetic graphs (20%)
(h) synthetic graphs (30%)
(i) synthetic graphs (40%)
(j) synthetic graphs (50%)
Figure 11. The DSM efficiency on edge-deletion-only real/synthetic graphs, compared with baseline methods.
@@ 8.2. Parameter Tuning
In this subsection, we tune parameters for our DSM approach over synthetic graphs.
The DSM Efficiency w.r.t. SPUR/SPAN Vector Dimension $\bm{d}$: Figure 8(a) illustrates the DSM performance by varying the dimension, $d$, of the cost-model-based SPUR/SPAN vector (in $o^{\prime}_{C}(v_{i})$ of Eq. (11)) from 1 to 4, where other parameters are set to their default values. With higher embedding dimension $d$, the pruning power of our proposed pruning strategies in higher dimensional space increases. However, the access of synopses with larger $d$ may also incur higher costs due to the ”dimensionality curse” (Berchtold et al., 1996). Thus, in this figure, for larger $d$, the wall clock time of DSM first decreases and then increases over all synthetic graphs. Nonetheless, the time cost remains low (i.e., 1.96 $\sim$ 6.05 $ms$) for different $d$ values.
The DSM Efficiency w.r.t. $\bm{\beta/\alpha}$ Ratio: Figure 8(b) varies the ratio, $\beta/\alpha$, from 10 to 100,000 for the optimized (or cost-model-based) vertex dominance embeddings $o^{\prime}(v_{i})$ (or $o^{\prime}_{C}(v_{i})$), where other parameters are set by default. In the figure, we can see that our DSM approaches are not very sensitive to the ratio $\beta/\alpha$. For different $\beta/\alpha$ ratios, the query cost remains low (i.e., 1.96 $\sim$ 3.75 $ms$).
The DSM Efficiency w.r.t. ##\## of Degree Groups, $\bm{m}$: Figure 8(c) reports the performance of our DSM approach, by varying the number, $m$, of degree groups from 1 to 5, where other parameters are set by default. In this figure, the time costs over $Syn\text{-}Uni$ and $Syn\text{-}Zipf$ first decrease and then increase when $m$ increases, and there are some fluctuations for $Syn\text{-}Gau$ (e.g., for $m=4$ or $5$). For all $m$ values, the time cost remains low (i.e., 1.94 $\sim$ 3.74 $ms$).
(a) DSM vs. $|\Sigma|$
(b) DSM vs. $avg\_deg(q)$
(c) DSM vs. $|V(q)|$
(d) DSM vs. $avg\_deg(G_{D})$
(e) DSM vs. $|V(G_{D})|$
Figure 12. The DSM efficiency w.r.t. parameters $|\Sigma|$, $avg\_deg(q)$, $|V(q)|$, $avg\_deg(G_{D})$, and $|V(G_{D})|$.
The DSM Efficiency w.r.t. ##\## of Cell Intervals on Each Dimension, $\bm{K}$: Figure 8(d) evaluates the effect of the number, $K$, of cell intervals on each dimension on the DSM performance, where $K$ varies from 1 to 10, and other parameters are set by default. When $K$ becomes larger, more vertices in synopsis cells can be pruned, however, more cells need to be accessed. Therefore, for DSM, with the increase of $K$, the time cost first decreases and then increases. Nonetheless, for different $K$ values, the query cost remains low (i.e., 1.96 $\sim$ 3.54 $ms$).
@@
The experimental results on real-world graphs are similar and thus omitted here. In subsequent experiments, we will set parameters $d=2$, $\beta/\alpha=1,000$, $m=3$, and $K=5$, and use the cost-model-based vertex embeddings $o^{\prime}_{C}(v_{i})$ (given by Eq. (11)).
8.3. The DSM Effectiveness Evaluation
In this subsection, we report the pruning power of our proposed pruning strategies (as discussed in Section 6.1) for dynamic subgraph matching queries over real/synthetic graphs.
The DSM Pruning Power on Real/Synthetic Graphs: Figure 9 shows the pruning power of our DSM approach over real/synthetic graphs with different embedding designs $o(v_{i})$, $o^{\prime}(v_{i})$, and $o^{\prime}_{C}(v_{i})$, where all parameters are set by default. In figures, we can see that for all real/synthetic graphs, the pruning power of our proposed vertex embedding designs can reach as high as $94.47\%\sim 99.58\%$ for real-world graphs and $85.99\%\sim 98.65\%$ for synthetic graphs, which confirms the effectiveness of our embedding-based pruning strategies. Our cost-model-based vertex embedding strategy $o^{\prime}_{C}(v_{i})$ always achieves the highest pruning power (i.e., $86.93\%\sim 99.58\%$).
8.4. The DSM Efficiency Evaluation
In this subsection, we report the efficiency of our DSM approaches over real/synthetic graphs.
The DSM Efficiency on Edge-Insertion-Only Real/Synthetic Graphs: Figure 10 compares the efficiency of our DSM approach with that of 5 state-of-the-art baseline methods by varying the insertion ratio from 10% to 50% over both real and synthetic graphs, where all other parameters are set to default values. In Figures 10(a) and 10(f) with the default 10% insertion ratio, we can see that our DSM approach outperforms baseline methods mostly by 1-2 orders of magnitude. Overall, for all graphs (even for $up$ with $3.77M$ vertices and $1.65M$ edge insertions), the time cost of our DSM approach remains low (i.e., $<$$3.91$ $sec$).
For other subfigures with different insertion ratios from 20% to 50%, we can see similar experimental results over both real and synthetic graphs. Although the total query time will increase as the number of inserted edges increases, our DSM approach can achieve performance that is up to 1-2 orders of magnitude than baseline methods. In summary, for all insertion ratios and graphs (even for $up$ with $3.77M$ vertices and $8.26M$ edge insertions), the time cost of our DSM approach remains low (i.e., $<4.37$ $sec$).
@@ 8.4. The DSM Efficiency Evaluation
From other subfigures with different deletion ratios 20% $\sim$ 50%, we can find similar experimental results over both real and synthetic graphs, where our DSM approach can always outperform the three baseline methods. For different edge deletion ratios and graph sizes (even for $up$ with $3.77M$ vertices and $8.26M$ edge deletions), the time cost of our DSM approach remains low (i.e., $<4.62$ $sec$).
To evaluate the robustness of our DSM approaches, in the sequel, we vary the parameter values on synthetic graphs (e.g., $|\Sigma|$, $avg\_deg(q)$, $|V(q)|$, $avg\_deg(G_{D})$, and $|V(G_{D})|$). To better illustrate the trends of curves, we omit the results of the baselines below.
The DSM Efficiency w.r.t. ##\## of Distinct Vertex Labels, $\bm{|\Sigma|}$: Figure 12(a) shows the wall clock time of our DSM approach, where $|\Sigma|$ varies from 5 to 25, and other parameters are set to default values. When the number, $|\Sigma|$, of distinct vertex labels increases, the pruning power also increases (i.e., with fewer candidate vertices). Moreover, the query cost is also affected by vertex label distributions. Overall, the DSM query cost remains low for different $|\Sigma|$ values (i.e., 1.83 $\sim$ 4.04 $ms$).
The DSM Efficiency w.r.t. Average Degree, $\bm{avg\_deg(q)}$, of the Query Graph $\bm{q}$: Figure 12(b) examines the DSM performance by varying the average degree, $avg\_deg(q)$, of the query graph $q$ from 2 to 4, where other parameters are set to default values. Higher degree $avg\_deg(q)$ of $q$ incurs higher pruning power of query vertices. Therefore, when $avg\_deg(q)$ increases, the DSM time cost decreases. For different $avg\_deg(q)$ values, the DSM query cost remains low (i.e., 1.58 $\sim$ 7.68 $ms$).
@@
The DSM Efficiency w.r.t. Avg. Degree, $\bm{avg\_deg(G_{D})}$, of Dynamic Graph $\bm{G_{D}}$: Figure 12(d) presents our DSM performance with different average degrees, $avg\_deg$ $(G_{D})$, of dynamic graph $G_{D}$, where $avg\_deg(G_{D})=3\sim 7$, and default values are used for other parameters. Intuitively, higher degree $avg\_deg(G_{D})$ in data graph $G_{D}$ incurs lower pruning power and more candidate vertices. Thus, when $avg\_deg(G_{D})$ becomes higher, the wall clock time also increases, especially for $Syn\text{-}Zipf$ (due to its skewed vertex label distribution). Nevertheless, the DSM query time remains small for different $avg\_deg(G_{D})$ values (i.e., 1.08 $\sim$ 5.82 $ms$).
(a) real-world graphs
(b) synthetic graphs
Figure 13. The comparison analysis of the DSM offline pre-computation and online query costs on real/synthetic graphs.
(a) real-world graphs
(b) synthetic graphs

### v3

@@ 7. Cost-Model-Guided Embedding Optimization
7.1. Cost Model for CSM via Vertex Embeddings
In this subsection, we will propose a cost model to evaluate the performance of our continuous subgraph matching (or the pruning power of using vertex dominance embeddings).
Specifically, we estimate the number, $Cost_{CSM}$, of candidate vertices $v_{i}$ (to retrieve and refine) whose embedding vectors $o^{\prime}(v_{i})$ are dominated by query embedding vector $o^{\prime}(q_{i})$ (given by Lemma 6.1):
(8) $Cost_{CSM}=\sum_{v_{i}\in V(G_{t})}\prod_{j=1}^{2d}Pr\{o^{\prime}(q_{i})[j]\leq o^{\prime}(v_{i})[j]\},$
@@ 7.1. Cost Model for CSM via Vertex Embeddings
$\Pr\{o^{\prime}(q_{i})[j]\leq o^{\prime}(v_{i})[j]\}$
[…]
(9) $\left.\quad\leq\frac{-(o^{\prime}(q_{i})[j]-\mu_{o^{\prime}(v_{i})[j]})}{\sigma_{o^{\prime}(v_{i})[j]}^{2}}\right\}.$
@@
(11) $Cost_{CSM}=|V(G_{t})|\cdot\prod_{j=1}^{2d}\Phi\left(\frac{\mu_{o^{\prime}(v_{i})[j]}-o^{\prime}(q_{i})[j]}{\sigma_{o^{\prime}(v_{i})[j]}^{2}}\right).$
7.2. Cost-Model-Based Vertex Embedding Design
Eq. (11) estimates the (worst-case) query cost for one query vertex $q_{i}$ during the continuous subgraph matching. Intuitively, our goal is to design vertex embeddings $o^{\prime}(\cdot)$ that minimizes the cost model $Cost_{CSM}$ (given in Eq. (11)).
Note that, $\Phi(\cdot)$ in Eq. (11) is a monotonically increasing function. Therefore, if we can minimize the term $\frac{\mu_{o^{\prime}(v_{i})[j]}-o^{\prime}(q_{i})[j]}{\sigma_{o^{\prime}(v_{i})[j]}^{2}}$ in Eq. (11), then low query cost $Cost_{CSM}$ can be achieved. In other words, guided by our proposed cost model (to minimize the query cost $Cost_{CSM}$), our target is to design/select a “good” distribution of vertex dominance embeddings $o^{\prime}(v_{i})$ with:
@@ 7.2. Cost-Model-Based Vertex Embedding Design
In this paper, unlike standard vertex dominance embeddings generated by Uniform random function $f_{1}(\cdot)$ (as discussed in Section 4.2), we choose to use a (seeded) Zipf random function to produce SPUR vectors $x_{i}$ in vertex embeddings $o^{\prime}(v_{i})$ (or in turn generate SPAN vectors $y_{i}=\sum x_{j}$). The $Zipf$ distribution exactly follows our target of finding a random variable with low mean and high variance, which can achieve low query cost, as guided by our proposed cost model $Cost_{CSM}$ (given in Eq. (8)). Please refer to experimental comparison of $Zipf$ with $Uniform$ in Figure 14(a) of Section 8. We would like to leave the interesting topic of studying other low-mean/high-variance distributions as our future work.
@@
8. Experimental Evaluation
8.1. Experimental Settings
To evaluate the performance of our DIVINE approach, we conduct experiments on a Ubuntu server equipped with an Intel Core i9-12900K CPU and 128GB memory. Our source code in C++ and real/synthetic graph data sets are available at URL: https://github.com/JamesWhiteSnow/DSM.
Real/Synthetic Graph Data Sets: We evaluate our DIVINE approach over both real and synthetic graphs.
Real-world graphs. We test six real-world graph data used by previous works (He and Singh, 2008; Shang et al., 2008; Zhao and Han, 2010; Lee et al., 2012; Sun et al., 2012; Han et al., 2013; Ren and Wang, 2015; Bi et al., 2016; Katsarou et al., 2017; Hamilton et al., 2017; Bhattarai et al., 2019; Han et al., 2019; Sun and Luo, 2020), which can be classified into three categories: i) biology networks (Yeast and HPRD); ii) bibliographical/social networks (Reddit, DBLP and Youtube), and; iii) citation networks (US Patents). Statistics of these real graphs are summarized in Table 2.
Synthetic graphs. We generate synthetic graphs via NetworkX (Hagberg and Conway, 2020), and produce small-world graphs following the Newman-Watts-Strogatz model (Watts and Strogatz, 1998). Parameter settings of synthetic graphs are depicted in Table 3. For each vertex $v_{i}$, we generate its label $l(v_{i})$ by randomly picking up an integer within $1,|\sum|]$, following the $Uniform$, GaussianGaussian, or $Zipf$ distribution. Accordingly, we obtain 3 types of graphs, denoted as $Syn\text{-}Uni$, $Syn\text{-}Gau$, and $Syn\text{-}Zipf$, respectively.
The insertion (or deletion) rate is defined as the ratio of the number of edge insertions (or deletions) to the total number of edges in the raw graph data. Following the literature (Choudhury et al., 2015; Kim et al., 2018; Min et al., 2021; Sun et al., 2022a), we set the insertion rate to 10% by default (i.e., initial graph $G_{0}$ contains 90% edges, and the remaining 10% edges arrive as the insertion stream). Similarly, for the case of edge deletions, we set the default deletion rate to 10% (i.e., initial graph $G_{0}$ contains all edges, and deletion operations of 10% edges arrive in the stream). Below, we will report the CSM performance with edge insertion only, unless specified otherwise (see the edge-deletion case in Figure LABEL:subfig:dsm_del_real_mix).
Query Graphs: Similar to previous works (Sun et al., 2012; Han et al., 2013; Ren and Wang, 2015; Bi et al., 2016; Katsarou et al., 2017; Archibald et al., 2019; Bhattarai et al., 2019; Han et al., 2019; Sun et al., 2022a), for each graph $G_{D}$, we randomly extract/sample 100 connected subgraphs as query graphs, where parameters of query graphs $q$ (e.g., $|V(q)|$ and avg_deg(q)avg\_deg(q)) are depicted in Table 3.
Baseline Methods: We compare our DIVINE approaches (using the cost-model-based vertex dominance embeddings $o^{\prime}_{C}(v_{i})$, as discussed in Section 7) with five representative baseline methods of dynamic subgraph matching and one GNN-based path embedding baseline of static subgraph matching as follows:
@@ 8.1. Experimental Settings
  6. (6)
    GNN-PE (Ye et al., 2024) retrains the GNN model for path embedding if the dominance relationship does not hold (due to dynamic graph updates).
We used the code of baseline methods from (Sun et al., 2022a; Ye et al., 2024), where GNN-PE is implemented by Python and C++, and the other methods are implemented in C++ for a fair comparison.
Table 3. Parameter settings.
Parameters Values
the dimension, $d$, of the SPUR/SPAN vector 1 2, 3, 4
the ratio, $\beta/\alpha$ 10, 100, 1,000, 10,000, 100,000
the number, $m$, of degree groups 1, 2, 3, 4, 5
the number, $K$, of cell intervals in each dimension of $Syn_{j}$ 1, 2, 5, 8, 10
the number, $|\Sigma|$, of distinct labels 5, 10, 15, 20, 25
the average degree, avg_deg(q)avg\_deg(q), of the query graph $q$ 2, 3, 4
the size, $|V(q)|$, of the query graph $q$ 5, 6, 8, 10, 12
the average degree, $avg\_deg(G_{D})$, of dynamic graph $G_{D}$ 3, 4, 5, 6, 7
the size, $|V(G_{D})|$, of the dynamic data graph $G_{D}$ 10K, 30K, 50K, 80K, 100K, 500K, 1M, 5M, 10M
Evaluation Metrics: In our experiments, we report the efficiency of our DIVINE approach and baseline methods, in terms of the wall clock time, which is the time cost of processing the CSM query over the entire graph update sequence (i.e., continuous subgraph matching answer set monitoring). In particular, this time cost includes the time of filtering, refinement, embedding updates, and synopsis updates for each CSM query. We also evaluate the pruning power of our embedding dominance pruning and MBR range pruning strategies (as mentioned in Section 6.1), which is defined as the percentage of vertices that can be ruled out by our pruning methods. For all experiments, we take an average of each metric over 100 runs (with 100 query graphs, respectively).
Table 3 depicts our parameter settings, where default parameter values are in bold. For each set of experiments, we vary the value of one parameter while setting other parameters to default values.
@@ 8.2. Parameter Tuning
We first tune the parameters for our DIVINE approach using synthetic graph data sets.
DIVINE vs. $d$] Refer to caption
DIVINE vs. $\beta/\alpha$] Refer to caption
DIVINE vs. $K$] Refer to caption
DIVINE vs. $m$] Refer to caption
DIVINE vs. embeddings] Refer to caption
(a) The DIVINE efficiency w.r.t parameters $d$, $\beta/\alpha$, $K$, $m$, and embedding design strategies.
The DIVINE Efficiency w.r.t. SPUR/SPAN Vector Dimension $\bm{d}$: Figure 8.2 illustrates the DIVINE performance by varying the dimension, $d$, of the cost-model-based SPUR/SPAN vector (in $o^{\prime}_{C}(v_{i})$ of Eq. (12)) from 1 to 4, where other parameters are set to their default values. With a higher embedding dimension $d$, the pruning power of our proposed pruning strategies in a higher-dimensional space increases. However, the access of synopses with larger $d$ may also incur higher costs due to the ”dimensionality curse” (Berchtold et al., 1996). Thus, in this figure, for larger $d$, the wall clock time of DIVINE first decreases and then increases over all synthetic graphs. Nonetheless, the time cost remains low (i.e., 1.96 $\sim$ 6.05 $ms$) for different $d$ values.
The DIVINE Efficiency w.r.t. $\bm{\beta/\alpha}$ Ratio: Figure 8.2 varies the ratio, $\beta/\alpha$, from 10 to 100,000 for the optimized (or cost-model-based) vertex dominance embeddings $o^{\prime}(v_{i})$ (or $o^{\prime}_{C}(v_{i})$), where other parameters are set by default. In the figure, we can see that our DIVINE approaches are not very sensitive to the ratio $\beta/\alpha$. For different $\beta/\alpha$ ratios, the query cost remains low (i.e., 1.96 $\sim$ 3.75 $ms$).
@@
The experimental results on real-world graphs are similar and thus omitted here.
In subsequent experiments, we will set parameters $d=2$, $\beta/\alpha=1,000$, $m=3$, and $K=5$, and use the best cost-model-based vertex embeddings $o^{\prime}_{C}(v_{i})$ (given by Eq. (12)).
8.3. The DIVINE Effectiveness Evaluation
In this subsection, we report the pruning power of our pruning strategies (as discussed in Section 6.1), as well as the query selectivity of our proposed embeddings, for continuous subgraph matching over real/synthetic graphs.
real-world graphs] Refer to caption
synthetic graphs] Refer to caption
(b) The DIVINE pruning power w.r.t. different design strategies of vertex dominance embeddings, compared with naive label filtering (NLF) and GNN-PE.
The DIVINE Pruning Power: Figure 14(b) shows the pruning power of our DIVINE approach (different embedding designs $o(v_{i})$, $o^{\prime}(v_{i})$, and $o^{\prime}_{C}(v_{i})$) over real/synthetic graphs, compared with that of naive label filtering (NLF) and GNN-PE (Ye et al., 2024), where all parameters are set by default. Specifically, NLF filters out candidate vertices with labels different from the query vertex, whereas GNN-PE rules out candidate vertices whose learned embedding vectors are not dominated by that of the query vertex. In the subfigures, we can see that our DIVINE approach even with the naive embedding design $o(v_{i})$ can significantly outperform both NLF and GNN-PE, over all the real/synthetic graphs.
Moreover, we can see that the DIVINE pruning power of our proposed vertex embedding designs can reach as high as $94.47\%\sim 99.58\%$ for real-world graphs and $85.99\%\sim 98.65\%$ for synthetic graphs, which confirms the effectiveness of our embedding-based pruning strategies. Our cost-model-based vertex embedding strategy $o^{\prime}_{C}(v_{i})$ always achieves the highest pruning power (i.e., $86.93\%\sim 99.58\%$). Thus, in the sequel, we will always use the cost-model-based vertex embeddings in our DIVINE approach.
real-world graphs] Refer to caption
synthetic graphs] Refer to caption
(c) The comparison of the selectivity distributions between DIVINE and NLF.
The DIVINE Selectivity Distribution: Figure 14(c) evaluates the query selectivity distribution of our DIVINE approach, compared with that of NFL, where default values are assigned to parameters. Specifically, we first obtain 800 query vertices from 100 query graphs (i.e., each graph with 8 query vertices), and then report the numbers of candidate vertices dominated by their corresponding 800 query embedding vectors (using DIVINE or NLF), respectively, in box plots (i.e., query selectivity distribution with MIN, MAX, medium, and lower/upper quantiles). From the subfigures, we can see that, compared with NLF, our DIVINE approach has significantly fewer candidate vertices for most query vertices, which indicates better selectivity of our DIVINE embedding design.
Figure 14. A visualization of vertex dominance embeddings.
The Visualization of Our Vertex Dominance Embedding: Figure 14 visualizes vertex dominance embeddings over the ye data set (gray points), where we plot 2D SPUR and SPAN vectors in the two leftmost figures. As a case study, given a query vertex $o^{\prime}(q_{i})$ (red star point), its dominating region $DR(o^{\prime}(q_{i}))$ contains a candidate vertex $v_{i}$ (green diamond) that matches with $q_{i}$. We can see that the subgraph relationship of star structures can be well-preserved if their vertex embeddings follow the dominance relationship.
8.4. The DIVINE Efficiency Evaluation
In this subsection, we report the efficiency of our DIVINE approaches over real/synthetic graphs.
real-world graphs ($Ins_{10\%}$)] Refer to caption
real-world graphs ($Ins_{20\%}$)] Refer to caption
real-world graphs ($Ins_{30\%}$)] Refer to caption
real-world graphs ($Ins_{40\%}$)] Refer to caption
real-world graphs ($Ins_{50\%}$)] Refer to caption
synthetic graphs ($Ins_{10\%}$)] Refer to caption
synthetic graphs ($Ins_{20\%}$)] Refer to caption
synthetic graphs ($Ins_{30\%}$)] Refer to caption
synthetic graphs ($Ins_{40\%}$)] Refer to caption
synthetic graphs ($Ins_{50\%}$)] Refer to caption